OpenStudy (anonymous):

PLEASE HELP!!! My assignment asked for me to draw a right triangle. I measured the sides and the hypotenuse measures 7.14, the opposite leg 3.5, and the adjacent side 7. I used the converse Pythagorean Theorem to verify that my triangle was a right triangle but the measures were not consistent. Why? I also need to write the sin, cosine, and tangent ratios for

OpenStudy (anonymous):

Your cos, tan, and sin measurements are all correct. 7 is wayy too big to be one of the legs if 7.25 is your hypotenuse. Do this instead, just plug in some numbers into a and b for the pythagorean theorem, and then when you find c make those measurements. It's okay, just a mistake in measuring :) for example, 2 in for one leg, and 3 in for another 2^2 + 3 ^2 = c^2 4 + 9 = c^2 square root of 13 = c